Vaughn Manufacturing used high-low data from June and July to determine its variable cost of $10 per unit. Additional information follows:

Month Units produced Total costs
June 3000 $41000
July 800 19000



If Vaughn’s produces 3000 units in August, how much is its total cost expected to be?

$17368

$41000

$30000

$52000
Answer #1

Total fixed costs would be=41000-(3000*$10)=$11,000

Hence total cost for 3000 units=Variable cost+Fixed costs

=11000+(10*3000)

which is equal to

=$41000

NOTE:Total fixed costs and variable cost per unit do not change with change in units